# Constraints from Entanglement Wedge Nesting for Holography at a Finite Cutoff

### v2

@@ 1.1 Motivation and Background
Concrete realizations for the surface/state correspondence can be found, apart from tensor networks which motivated the correspondence, for example in the study of $T\bar{T}$-deformations of holographic CFTs Zamolodchikov:2004ce ; Jiang2019LecturesOS ; Bonelli:2018kik ; Dubovsky:2017cnj ; Grieninger:2019zts ; Gross:2019ach ; Cavaglia:2016oda ; Pant:2024eno . It has been proposed that a $T\bar{T}$ deformation of a holographic CFT should be understood as “moving the boundary into the bulk” McGough:2016lol ; Guica:2019nzm ; Lewkowycz:2019xse . Furthermore, the appearance of cutoff surfaces have also played a key role in double holography Takayanagi:2011zk ; Almheiri:2019hni ; Geng:2020fxl ; Geng:2020qvw ; Chen:2020uac ; Chen:2020hmv ; Hernandez:2020nem ; Grimaldi:2022suv ; Geng:2021wcq ; Geng:2021iyq ; Neuenfeld2021DoubleHA ; Neuenfeld:2021wbl ; Omiya:2021olc ; Geng:2023iqd ; Geng:2023qwm , where a convex cutoff surface in the form of an end-of-the-world (ETW) brane appears. Doubly-holographic models with AdS branes start by introducing the dual of a holographic $d$-dimensional boundary BCFT in terms of a $d+1$-dimensional bulk AdS spacetime. [^1] BCFTs are conformal field theories living on a space with boundary at which conformal boundary conditions are imposed. However, at least for a BCFT2 one needs to introduce boundary degrees of freedom which break conformal invariance in order to obtain interesting models, e.g. by coupling an SYK model to the BCFT boundary. We will will thus use the term BCFT more loosely to also account for situations like this. The boundary of the BCFT is continued into the bulk by an ETW brane Takayanagi:2011zk which removes part of the naive bulk spacetime dual to a CFT without a boundary.
@@
where $\chi(U)$ is the the minimal area extremal surface homologous to the boundary subregion $U$. The surface $\chi(U)$ is known as the HRRT surface or RT surface for short.
The surface/state correspondence proposes that in the case of a finite cutoff the von Neumann entropy of a reduced density matrix obtained from |$\ket{\phi(\Sigma)}$ can be computed from co-dimension two extremal surfaces that are anchored at the boundary of subregions of (a slice of) the spacetime boundary $\Sigma$. A basic consistency check of this proposal with holographic subregion entropy computations has e.g. been performed in the context of $T\bar{T}$ in Donnelly:2018bef . Also in double holography, there is evidence that subregion entropies on the brane can be computed using the RT formula Chen:2020uac . However, it is still unknown under which conditions the proposal of the surface/state correspondence is fully consistent. While it is not even clear what “consistency” entails in detail, strong subadditivity (SSA) of holographic entanglement entropy Headrick:2007km ; Wall:2012uf should certainly be a necessary condition if we want the holographic entanglement entropy to behave like an entropy in the quantum information theoretic sense.
@@
We will assume that the RT formula in its simplest formulation, Eq. (1), holds [^3] The authors of Bousso:2022hlz proposed the notion of generalized entanglement wedges for gravitational theories. Their proposal appears to only work for subregions of a time-reflection symmetric Cauchy slice and thus is not applicable to our configurations. . Given this starting point, the RT surface which computes $S(A\cup B)$ for two disconnected, spacelike subregions $A$ and $B$ consists of two disconnected pieces which can be in one of two phases. Either, each connected piece can be homologous to one of $A$ and $B$. In this case, which we will call the disconnected phase, the mutual information $S(A)+S(B)-S(A\cup B)$ vanishes. The second phase has the RT surface connect the subregions $A$ and $B$ to each other. Consequently, the mutual information does not vanish. We will call this the connected phase. We will then construct the entanglement wedges $\mathcal{W}_{E}$ for the regions $A$, $B$, and $A\cup B$. As we will argue below, while the definitions of $\mathcal{W}_{E}(A)$ and $\mathcal{W}_{E}(B)$ are straight forward and agree with the intersection of entanglement wedge in the full geometry cut off by the brane, the definition of the connected entanglement wedge is more subtle and requires us to be careful about the location and direction at which the RT surface ends on the cutoff surface.
@@
It is well known Lewkowycz:2019xse ; Geng:2019ruz ; Grado-White:2020wlb ; Neuenfeld:2023svs that naively extending in HRT formula for cutoff surfaces can yield situation where SSA and EWN is violated. This motivated the authors of Grado-White:2020wlb to propose a modification of the RT formula for cutoff surfaces, called restricted maximin. The main idea of their proposal is to only consider bulk extremal surfaces which lie on a Cauchy slice that intersects $U$, which guarantees SSA and EWN. However, as we will argue below, applying their construction to a configuration of disjoint subregions $A$ and $B$ which violate our bounds will produce RT surfaces which disagree with naive expectations. For example, given a static geometry and subregions $A$ and $B$ on constant but different time-slices, the two disjoint components of the RT surface in the disconnected phase will have a non-trivial profile in time. Since this happens whenever $\mathcal{W}_{E}(A)$ and $\mathcal{W}_{E}(B)$ are timelike separated, this suggests that a deviation between the naive RT surfaces defined in this paper and restricted maximin surfaces indicates that $A$ and $B$ are not independent, possibly due to the non-local nature of the theory at the cutoff.
The issue of defining the subregion, $B$, on the cutoff/brane can be subtle depending on the types of boundary condition (BC) we impose on the brane. When discussing the brane/cutoff there appear to be at least two distinct ways in which we might interpret the physics on the brane which depends on whether we choose Dirichlet or Neumann BCs on the brane Geng:2023qwm ; Deng:2023pjs ; Geng:2024xpj . When using Neumann BCs the way the brane is embedded in the higher dimensional AdS geometry is determined dynamically by analyzing equations of motion of the following Lorentzian action [^4] Note in our setup in Poincare AdS the boundary of the bulk cutoff by the brane consists of the conformal boundary as well as the ETW brane so we should also include a Gibbons-Hawking term of the form $\int_{\text{bdry}}d^{d}y\sqrt{-h_{\text{bdry}}}K_{\text{bdry}}$. For simplicity we omit explicitly writing it this in the action and focus on the brane and bulk terms. Takayanagi:2011zk
@@
which implies the brane/cutoff is a co-dimension one surface in the bulk of constant extrinsic curvature characterized by a tension $T$ and usually one can consider dynamical gravity on the brane as is often done in double holography Takayanagi:2011zk ; Almheiri:2019hni ; Geng:2020fxl ; Geng:2020qvw ; Chen:2020uac ; Chen:2020hmv ; Hernandez:2020nem ; Grimaldi:2022suv ; Neuenfeld2021DoubleHA ; Neuenfeld:2021wbl ; Omiya:2021olc ; Geng:2023iqd ; Geng:2023qwm and brane-world gravity Karch:2000ct ; Geng:2023qwm . This naively poses the questions of how to define cutoff subregions in a diffeomorphism invariant. One possible way to address this by going back to the subregion $A$ which lies on the non-gravitating conformal boundary. Then we consider RT surfaces associated to $A$ which extend to the brane. In the case of Neumann BCs the location the RT surfaces end on the brane is dynamically determined and this in turn can be used to associate a subregion $B$ on the brane given some $A$ on the conformal boundary (i.e. $B$ is not independently chosen from $A$). Since we are considering dynamical gravity on the brane one can change the gravitational dynamics on the brane by adding terms to Eq. (2) (e.g, higher derivative corrections, DGP gravity, etc. on the brane Dvali:2000hr ; Chen:2020hmv ; Chen:2020uac ; Geng:2022tfc ; Geng:2022slq ; Geng:2023qwm ; Lee:2022efh ) these would generally affect where the RT surface ends on the brane and thereby shift the relative location of $A$ and $B$. One might speculate that perhaps one might end up in configurations that violate EWN, in this case the constraints we find may be regarded as providing constraints to gravity on the brane. On the other hand, imposing Dirichlet boundary conditions on the brane one can no longer view the embedding of the brane in AdS as arising from the equation of motion given in Eq. (4). In this case we simply view the brane as a cutoff surface of constant extrinsic curvature where all bulk fields satisfy Dirichlet boundary conditions on the brane, this is more akin to an interpretation of a cutoff in the $T\bar{T}$ context where gravity is effectively turned off on the brane. In such a case we can fix some gauge and naively choose spacelike separated subregions $A$ and $B$ “independently” of each other. In this case what we constrain through EWN is not gravity on the brane as we did in the Neumann case but rather the relative location of $A$ and $B$. The central point here is that regardless of the interpretation we place on the physics of the brane our results simply rely on having two subregions; $A$ on a conformal boundary and $B$ on the brane/cutoff along with a prescription to associate entanglement wedges to these subregions. With this we can derive constraints from EWN and such an investigation on its own will yield interesting results which we summarize in the following section.
@@ 1.2 Overview
In Section 4 we study EWN in a two-sided planar BTZ black hole background whose right exterior is cutoff by an ETW brane. We introduce the intervals $A_{\text{Br}}$ and $A_{\text{Bdry}}$ on the brane in the right exterior and on the boundary in the left exterior, respectively. We discuss the RT surfaces that will be homologous to the subregions $A_{\text{Br}}$, $A_{\text{Bdry}}$, and $A_{\text{Bdry}}\cup A_{\text{Br}}$ and use them to define the entanglement wedges $\mathcal{W}_{E}(A_{\text{Bdry}})$, $\mathcal{W}_{E}(A_{\text{Br}})$, and $\mathcal{W}_{E}(A_{\text{Bdry}}\cup A_{\text{Br}})$ respectively. Using similar methods employed in the first part of the paper in AdS3 we can write down the condition that needs to be placed on the intervals $A_{\text{Br}}$ and $A_{\text{Bdry}}$ so that $\mathcal{W}_{E}(A_{\text{Br}})\subset\mathcal{W}_{E}(A_{\text{Bdry}}\cup A_{\text{Br}})$ in the connected phase. We find that in the limit where the brane and boundary interval lengths diverge the condition has a simple geometric interpretation in the bulk. This interpretation is that requiring EWN in the connected phase prevents the connected RT surfaces from touching the black hole “singularities” at Kruskal time $\tau=\pm\pi/2$. This provides an interesting result in the context of Lee:2022efh which also analyzed the same BTZ black hole setup to constrain DGP couplings (more specifically the gravitational coupling set by the value of a JT dilaton) on the brane. The result we obtain serves as another reason beyond the ones discussed in Lee:2022efh to ignore RT surfaces that cross the “singularities” at $\tau=\pm\pi/2$.
@@ 2 Constructing Entanglement Wedges in AdS3 cutoff by ETW Brane
Figure 1: This plot showcases an example configuration of intervals $A$ and $B$, with $a_{1}=2,a_{2}=6,b_{1}=2,b_{2}=8$. The ETW brane with $\theta_{0}=\pi/4$ cuts off the shaded part of the spacetime. We have set $|t_{b}-t_{a}|=0.97(2+\sqrt{2})\approx 3.3$. The blue curve is $\chi_{\text{dis}}(A)$ and is anchored to the conformal boundary subregion $A$ and lies on the constant time slice $t=t_{a}$. The constant-time, solid yellow curve is $\chi_{\text{dis}}(B)$ and is anchored to the subregion $B$ on the brane. The two green curves show a projection of $\chi_{\text{con}}(A\cup B)=\chi_{1}\cup\chi_{2}$ and connect $A$ and $B$. For our our particular choice of parameters we have $|k_{1}|=0.97$ and $|k_{2}|=0.97\left[\frac{2-\sqrt{2}}{2}\right]\approx 0.28$ which implies that both green line segments are ellipses in the $xz$-plane. Note that all curves defined can be extended behind the brane (in the shaded region), but our constructions will only rely on the portion of these line segments strictly in unshaded part of the spacetime.
@@ 2.1 AdS3 Poincare Patch with an ETW Brane
where $\theta_{0}\in(0,\pi/2]$. As we already discussed in the introduction, when one imposes Neumann boundary conditions on the brane, the brane is a surface which will satisfy Eq. (4) with a tension given by $TL=\sin\theta_{0}$. When the tension is zero the ETW brane cuts off half of the AdS space and when $T=L^{-1}$ the ETW brane coincides with the conformal boundary. The limit of $T\lesssim\frac{1}{L}$, i.e., the limit in which the brane approaches the “would be conformal boundary” is referred to as the critical limit. The critical limit is physically relevant in the discussion of gravity induced on the brane, particularly in higher dimensions. It is a limit in which bulk gravitons become localized on the brane and one can consider a conventional lower dimensional gravity theory on the world volume of the brane Karch:2000ct ; Geng:2020qvw ; Neuenfeld:2021wbl . In the three-dimensional case discussed here, a theory of dilaton gravity gets induced on the brane, and the critical limit is the limit in which the cutoff scale of the brane theory becomes small Chen:2020uac ; Geng:2022slq ; Geng:2022tfc ; Neuenfeld:2023svs . In the case of Dirichlet BCs we simply interpret the brane as some cutoff surface of constant extrinsic curvature $K=2L^{-1}\sin\theta_{0}$, in this perspective there is no dynamical gravity on the brane [^5] With Dirichlet BCs the brane is more similar to a cutoff surface discussed in the $T\bar{T}$ context, whereas Neumann BCs is more appropriate in the discussion of islands in double holography and dynamical gravity on the brane.. In either perspective the key parameter which dictates how the brane/cutoff is embedded in the bulk is given by $\theta_{0}$ and all our results will be formulated in terms of $\theta_{0}$.

### v3

@@ 1.1 Motivation and Background
Concrete realizations for the surface/state correspondence can be found, apart from tensor networks which motivated the correspondence, for example in the study of $T\bar{T}$-deformations of holographic CFTs Zamolodchikov:2004ce ; Jiang2019LecturesOS ; Bonelli:2018kik ; Dubovsky:2017cnj ; Grieninger:2019zts ; Gross:2019ach ; Cavaglia:2016oda ; Pant:2024eno . It has been proposed that a $T\bar{T}$ deformation of a holographic CFT should be understood as “moving the boundary into the bulk” McGough:2016lol ; Guica:2019nzm ; Lewkowycz:2019xse . Furthermore, the appearance of cutoff surfaces have also played a key role in double holography Takayanagi:2011zk ; Almheiri:2019hni ; Geng:2020fxl ; Geng:2020qvw ; Chen:2020uac ; Chen:2020hmv ; Hernandez:2020nem ; Grimaldi:2022suv ; Geng:2021wcq ; Geng:2021iyq ; Neuenfeld2021DoubleHA ; Neuenfeld:2021wbl ; Omiya:2021olc ; Geng:2023iqd ; Geng:2023qwm , where a convex cutoff surface in the form of an end-of-the-world (ETW) brane appears. Doubly-holographic models with AdS branes start by introducing the dual of a holographic $d$-dimensional boundary BCFT in terms of a $d+1$-dimensional bulk AdS spacetime. [^1] BCFTs are conformal field theories living on a space with boundary at which conformal boundary conditions are imposed. However, at least for a BCFT2 one needs to introduce boundary degrees of freedom which break conformal invariance in order to obtain interesting models, e.g. by coupling an SYK model to the BCFT boundary. We will thus use the term BCFT more loosely to also account for situations like this. The boundary of the BCFT is continued into the bulk by an ETW brane Takayanagi:2011zk which removes part of the naive bulk spacetime dual to a CFT without a boundary.
@@
where $\chi(U)$ is the minimal area extremal surface homologous to the boundary subregion $U$. The surface $\chi(U)$ is known as the HRRT surface or RT surface for short.
The surface/state correspondence proposes that in the case of a finite cutoff the von Neumann entropy of a reduced density matrix obtained from |$\ket{\phi(\Sigma)}$ can be computed from co-dimension two extremal surfaces that are anchored at the boundary of subregions of (a slice of) the spacetime boundary $\Sigma$. A basic consistency check of this proposal with holographic subregion entropy computations has e.g. been performed in the context of $T\bar{T}$ in Donnelly:2018bef . Also, in double holography, there is evidence that subregion entropies on the brane can be computed using the RT formula Chen:2020uac . However, it is still unknown under which conditions the proposal of the surface/state correspondence is fully consistent. While it is not even clear what “consistency” entails in detail, strong subadditivity (SSA) of holographic entanglement entropy Headrick:2007km ; Wall:2012uf should certainly be a necessary condition if we want the holographic entanglement entropy to behave like an entropy in the quantum information theoretic sense.
@@
We will assume that the RT formula in its simplest formulation, Eq. (1), holds. [^3] The authors of Bousso:2022hlz proposed the notion of generalized entanglement wedges for gravitational theories. Their proposal appears to only work for subregions of a time-reflection symmetric Cauchy slice and thus is not applicable to our configurations. Given this starting point, the RT surface which computes $S(A\cup B)$ for two disconnected, spacelike subregions $A$ and $B$ consists of two disconnected pieces which can be in one of two phases. Either, each connected piece can be homologous to one of $A$ and $B$. In this case, which we will call the disconnected phase, the mutual information $S(A)+S(B)-S(A\cup B)$ vanishes. The second phase has the RT surface connect the subregions $A$ and $B$ to each other. Consequently, the mutual information does not vanish. We will call this the connected phase. We will then construct the entanglement wedges $\mathcal{W}_{E}$ for the regions $A$, $B$, and $A\cup B$. As we will argue below, while the definitions of $\mathcal{W}_{E}(A)$ and $\mathcal{W}_{E}(B)$ are straightforward and agree with the intersection of entanglement wedge in the full geometry cut off by the brane, the definition of the connected entanglement wedge is more subtle and requires us to be careful about the location and direction at which the RT surface ends on the cutoff surface.
@@
It is well known Lewkowycz:2019xse ; Geng:2019ruz ; Grado-White:2020wlb ; Neuenfeld:2023svs that naively extending the HRT formula for cutoff surfaces can yield situations where SSA and EWN is violated. This motivated the authors of Grado-White:2020wlb to propose a modification of the RT formula for cutoff surfaces, called restricted maximin. The main idea of their proposal is to only consider bulk extremal surfaces which lie on a Cauchy slice that intersects $U$, which guarantees SSA and EWN. However, as we will argue below, applying their construction to a configuration of disjoint subregions $A$ and $B$ which violate our bounds will produce RT surfaces which disagree with naive expectations. For example, given a static geometry and subregions $A$ and $B$ on constant but different time-slices, the two disjoint components of the RT surface in the disconnected phase will have a non-trivial profile in time. Since this happens whenever $\mathcal{W}_{E}(A)$ and $\mathcal{W}_{E}(B)$ are timelike separated, this suggests that a deviation between the naive RT surfaces defined in this paper and restricted maximin surfaces indicates that $A$ and $B$ are not independent, possibly due to the non-local nature of the theory at the cutoff.
The issue of defining the subregion, $B$, on the brane/cutoff can be subtle depending on the types of boundary condition (BC) we impose on the brane. When discussing the brane/cutoff there appear to be at least two distinct ways in which we might interpret the physics on the brane which depends on whether we choose Dirichlet or Neumann BCs on the brane Geng:2023qwm ; Deng:2023pjs ; Geng:2024xpj . When using Neumann BCs the way the brane is embedded in the higher dimensional AdS geometry is determined dynamically by analyzing equations of motion of the following Lorentzian action [^4] Note in our setup in Poincare AdS the boundary of the bulk cutoff by the brane consists of the conformal boundary as well as the ETW brane so we should also include a Gibbons-Hawking term of the form $\int_{\text{bdry}}d^{d}y\sqrt{-h_{\text{bdry}}}K_{\text{bdry}}$. For simplicity we omit explicitly writing it this in the action and focus on the brane and bulk terms. Takayanagi:2011zk
@@
which implies the brane/cutoff is a co-dimension one surface in the bulk of constant extrinsic curvature characterized by a tension $T$. Usually one can also consider dynamical gravity on the brane as is often done in double holography Takayanagi:2011zk ; Almheiri:2019hni ; Geng:2020fxl ; Geng:2020qvw ; Chen:2020uac ; Chen:2020hmv ; Hernandez:2020nem ; Grimaldi:2022suv ; Neuenfeld2021DoubleHA ; Neuenfeld:2021wbl ; Omiya:2021olc ; Geng:2023iqd ; Geng:2023qwm and brane-world gravity Karch:2000ct ; Geng:2023qwm . This naively poses the questions of how to define cutoff subregions in a diffeomorphism invariant way. One possible way to address this by going back to the subregion $A$ which lies on the non-gravitating conformal boundary. Then we consider RT surfaces associated to $A$ which extend to the brane. In the case of Neumann BCs the location the RT surfaces end on the brane is dynamically determined and this in turn can be used to associate a subregion $B$ on the brane given some $A$ on the conformal boundary (i.e. $B$ is not independently chosen from $A$). Since we are considering dynamical gravity on the brane one can change the gravitational dynamics on the brane by adding terms to Eq. (2) (e.g, higher derivative corrections, DGP gravity, etc. on the brane Dvali:2000hr ; Chen:2020hmv ; Chen:2020uac ; Geng:2022tfc ; Geng:2022slq ; Geng:2023qwm ; Lee:2022efh ) these would generally affect where the RT surface ends on the brane and thereby shift the relative location of $A$ and $B$. One might speculate that it is possible to end up in configurations that violate EWN, in this case the constraints we find may be regarded as providing constraints to gravity on the brane. On the other hand, imposing Dirichlet boundary conditions on the brane one can no longer view the embedding of the brane in AdS as arising from the equation of motion given in Eq. (4). In this case we simply view the brane as a cutoff surface of constant extrinsic curvature where all bulk fields satisfy Dirichlet boundary conditions on the brane, this is more akin to an interpretation of a cutoff in the $T\bar{T}$ context where gravity is effectively turned off on the brane. In such a case we can fix some gauge and naively choose spacelike separated subregions $A$ and $B$ “independently” of each other. In this case what we constrain through EWN is not gravity on the brane as we did in the Neumann case but rather the relative location of $A$ and $B$. The central point here is that regardless of the interpretation we place on the physics of the brane our results simply rely on having two subregions; $A$ on a conformal boundary and $B$ on the brane/cutoff along with a prescription to associate entanglement wedges to these subregions. With this we can derive constraints from EWN and such an investigation on its own will yield interesting results which we summarize in the following section.
@@ 1.2 Overview
In Section 4 we study EWN in a two-sided planar BTZ black hole background whose right exterior is cutoff by an ETW brane. We introduce the intervals $A_{\text{Br}}$ and $A_{\text{Bdry}}$ on the brane in the right exterior and on the boundary in the left exterior, respectively. We discuss the RT surfaces that will be homologous to the subregions $A_{\text{Br}}$, $A_{\text{Bdry}}$, and $A_{\text{Bdry}}\cup A_{\text{Br}}$ and use them to define the entanglement wedges $\mathcal{W}_{E}(A_{\text{Bdry}})$, $\mathcal{W}_{E}(A_{\text{Br}})$, and $\mathcal{W}_{E}(A_{\text{Bdry}}\cup A_{\text{Br}})$ respectively. Using similar methods employed in the first part of the paper we can write down the condition that needs to be placed on the intervals $A_{\text{Br}}$ and $A_{\text{Bdry}}$ so that $\mathcal{W}_{E}(A_{\text{Br}})\subset\mathcal{W}_{E}(A_{\text{Bdry}}\cup A_{\text{Br}})$ in the connected phase. We find that in the limit where the brane and boundary interval lengths diverge the condition has a simple geometric interpretation in the bulk. This interpretation is that requiring EWN in the connected phase prevents the connected RT surfaces from touching the black hole “singularities” at Kruskal time $\tau=\pm\pi/2$. This provides an interesting result in the context of Lee:2022efh which also analyzed the same BTZ black hole setup to constrain DGP couplings (more specifically the gravitational coupling set by the value of a JT dilaton) on the brane. The result we obtain serves as another reason beyond the ones discussed in Lee:2022efh to ignore RT surfaces that cross the “singularities” at $\tau=\pm\pi/2$.
@@ 2 Constructing Entanglement Wedges in AdS3 cutoff by ETW Brane
Figure 1: This plot showcases an example configuration of intervals $A$ and $B$, with $a_{1}=2,a_{2}=6,b_{1}=2,b_{2}=8$. The ETW brane with $\theta_{0}=\pi/4$ cuts off the shaded part of the spacetime. We have set $|t_{b}-t_{a}|=0.97(2+\sqrt{2})\approx 3.3$. The blue curve is $\chi_{\text{dis}}(A)$ and is anchored to the conformal boundary subregion $A$ and lies on the constant time slice $t=t_{a}$. The constant-time, solid yellow curve is $\chi_{\text{dis}}(B)$ and is anchored to the subregion $B$ on the brane. The two green curves show a projection of $\chi_{\text{con}}(A\cup B)=\chi_{1}\cup\chi_{2}$ and connect $A$ and $B$. For our particular choice of parameters we have $|k_{1}|=0.97$ and $|k_{2}|=0.97\left[\frac{2-\sqrt{2}}{2}\right]\approx 0.28$ which implies that both green line segments are ellipses in the $xz$-plane. Note that all curves defined can be extended behind the brane (in the shaded region), but our constructions will only rely on the portion of these line segments strictly in unshaded part of the spacetime.
@@ 2.1 AdS3 Poincare Patch with an ETW Brane
where $\theta_{0}\in(0,\pi/2]$. As we already discussed in the introduction, when one imposes Neumann boundary conditions on the brane, the brane is a surface which will satisfy Eq. (4) with a tension given by $TL=\sin\theta_{0}$. When the tension is zero the ETW brane cuts off half of the AdS space and when $T=L^{-1}$ the ETW brane coincides with the conformal boundary. The limit of $T\lesssim\frac{1}{L}$, i.e., the limit in which the brane approaches the “would be conformal boundary” is referred to as the critical limit. The critical limit is physically relevant in the discussion of gravity induced on the brane, particularly in higher dimensions. It is a limit in which bulk gravitons become localized on the brane and one can consider a conventional lower dimensional gravity theory on the world volume of the brane Karch:2000ct ; Geng:2020qvw ; Neuenfeld:2021wbl . In the three-dimensional case discussed here, a theory of dilaton gravity gets induced on the brane, and the critical limit is the limit in which the cutoff scale of the brane theory becomes small Chen:2020uac ; Geng:2022slq ; Geng:2022tfc ; Neuenfeld:2023svs . In the case of Dirichlet BCs we simply interpret the brane as some cutoff surface of constant extrinsic curvature $K=2L^{-1}\sin\theta_{0}$, in this perspective there is no dynamical gravity on the brane. [^5] With Dirichlet BCs the brane is more similar to a cutoff surface discussed in the $T\bar{T}$ context, whereas Neumann BCs is more appropriate in the discussion of islands in double holography and dynamical gravity on the brane. In either perspective the key parameter which dictates how the brane/cutoff is embedded in the bulk is given by $\theta_{0}$ and all our results will be formulated in terms of $\theta_{0}$.